# Hybrid Action Based Reinforcement Learning for Multi-Objective Compatible Autonomous Driving

### v2

@@ Abstract
Reinforcement Learning (RL) has shown excellent performance in solving decision-making and control problems of autonomous driving, which is increasingly applied in diverse driving scenarios. However, driving is a multi-attribute problem, leading to challenges in achieving multi-objective compatibility for current RL methods, especially in both policy execution and policy iteration. On the one hand, the common action space structure with single action type limits driving flexibility or results in large behavior fluctuations during policy execution. On the other hand, the multi-attribute weighted single reward function result in the agent’s disproportionate attention to certain objectives during policy iterations. To this end, we propose a Multi-objective Ensemble-Critic reinforcement learning method with Hybrid Parametrized Action for multi-objective compatible autonomous driving. Specifically, a parameterized action space is constructed to generate hybrid driving actions, combining both abstract guidance and concrete control commands. A multi-objective critics architecture is constructed considering multiple attribute rewards, to ensure simultaneously focusing on different driving objectives. Additionally, uncertainty-based exploration strategy is introduced to help the agent faster approach viable driving policy. The experimental results in both the simulated traffic environment and the HighD dataset demonstrate that our method can achieve multi-objective compatible autonomous driving in terms of driving efficiency, action consistency, and safety. It enhances the general performance of the driving while significantly increasing training efficiency.
@@ I Introduction
However, current RL methods still have several limitations in terms of compatibility with objectives such as driving efficiency, action consistency, and safety [8]. This does not satisfy the needs of multi-attribute driving tasks [9]. There are two main reasons: i) The connection between action space structure and actual driving behavior during policy execution is overlooked; ii) With a large state space and multiple attributes strongly coupled in open driving environments, the agent has difficulty achieving efficient policy iteration while being multi-objective compatible. These two reasons cause limitations in policy performance.
@@
For policy execution, directly utilizing a single-type action space to generate abstract or concrete driving behavior makes it challenging for the RL agent to both maintain flexibility and reduce driving behavior fluctuations. Specifically, a common approach involves having the agent generate discrete abstract driving goals, such as semantic decisions [10] or target points that guide path planning [11]. However, since the agent does not directly control the vehicle’s movement, its ability to influence driving flexibility is limited. Recently, it has become a popular approach for agent to directly output concrete control commands [12]. However, these commands, generated directly by the network, are prone to frequent fluctuations and abrupt responses to dynamic environmental changes.
In terms of policy iteration, it is mainly carried out through policy evaluation and policy exploration. For evaluation, a single reward function may not be compatible with all driving objectives, potentially causing the policy converge to local optimal. Specifically, when multiple attributes of an AD task are weighted into a single reward function, the agent may allocate disproportionate attention to certain attributes during training [8]. In an attempt to maximize rewards in certain states, some attributes may be neglected, leading to an inaccurate estimation of the value of those states. This may result in the driving performance being incompatible with the expectation of multiple objectives, such as becoming overly aggressive to maximize speed or excessively conservative to ensure safety. For exploration, performing random policy exploration based only on evaluated values lacks the orientation, causing the agent unable to actively explore unknown regions to discover potentially viable policies [13]. This random mechanism will lead agent to collect numerous repetitive exploration experiences that contribute little to the policy iterative update, resulting in inefficient policy convergence.
@@
To address the above issues, this paper proposes an Multi-objective Ensemble-Critic reinforcement learning method with Hybrid Parametrized Action space (HPA-MoEC) for multi-objective compatible autonomous driving. Our hybrid parametrized action space includes a discrete action set and corresponding continuous parameters, which generate driving actions that combine both the abstract guidance and concrete control commands. Building on this, we define multiple reward functions to decouple the attributes, with each reward function guiding one critic (i.e., value network) to focus on different objectives, cooperatively helping the actor (i.e., policy network) learn multi-objective compatible policies. Meanwhile, we introduce uncertainty to guide the agent in exploring viable driving policies, enabling more efficient exploration in unknown environments. The main contributions of our method are summarized as follows:
@@
    A hybrid parameterized action-based RL framework is proposed, which combines finer-grained guidance and control commands. The hybrid action space has consistently optimal discrete actions and their corresponding continuous action parameters. They further together generate abstract guidance and concrete control command outputs. Our method effectively achieves higher driving flexibility and smaller behavior fluctuations, ensuring the compatible in driving efficiency and action consistency.
@@
    A multi-objective compatible policy evaluation module is established, where critics focus on different objectives based on distinct reward functions. Given the safety-critical nature of AD, we set two driving objectives and evaluate them with two critics to verify the effectiveness of our design. One focuses on general performance, including interactivity, and the other on safety. Our design promotes multi-objective compatibility, with improvements in both general performance and safety demonstrated in the experimental results.
@@
    An epistemic uncertainty-based exploration strategy is designed for hybrid action, employing an ensemble of critics. By dynamically adjusting the direction and extent of exploration based on uncertainty and its changing trends, the agent is encouraged to faster explore regions of higher uncertainty for potentially viable policies. Our exploration strategy improves the learning efficiency of viable multi-objective compatible policies.
@@ II Related Works
The AD task involves making complex sequential decisions in a dynamic environment and can therefore be modeled as Markov Decision Processes (MDPs) [14]. The MDP is commonly represented as a tuple $<{\cal S},{\cal A},{\cal R},{\cal T},\gamma>$ , where ${\cal S}$ is the state space, ${\cal A}$ is the action space, ${\cal R}$ is the reward function, ${\cal T}$ is the transition function, and $\gamma$ is the discount factor. At time $t$, the RL agent selects action $a_{t}\in{\cal A}$ based on state $s_{t}\in{\cal S}$, then receives reward $r_{t}\in{\cal R}$ from the environment and transitions to state $s_{t+1}$ according to ${\cal T}$. The goal of the agent is to find an optimal policy through trial-and-error to maximize the expected reward. Next, we will introduce three aspects of related work on our method:
II-A Action Space Structure
Many current RL methods use a single action type to control vehicle driving, which fail to be compatible with high flexibility and small behavior fluctuations. On one hand, some studies use a discrete action space to generate abstract behavior decisions, offering long-term targets that indirectly guide vehicle control. Specifically, [15, 16] use DQN and its improved versions to generate semantic lateral actions, such as left or right lane changes. Additionally, [17, 18] introduces longitudinal discrete acceleration and deceleration actions. To provide clearer guidance, some studies select from a discrete set of trajectories [19, 20] or directly generate the positions and desired speeds of target points [21]. However, these methods often reduce the alignment between agent outputs and driving behavior, as they rely on integration with a basic controller for vehicle control, which limits flexibility. On the other hand, some studies [22, 23] directly generate steering angles laterally and accelerations longitudinally from a continuous action space, aiming to enhance flexibility. However, fluctuations in the network’s output can cause frequent changes in steering angle and acceleration commands [24]. In scenarios with dedicated lanes, lateral fluctuations caused by steering angle variations will result in unpredictable paths. The experimental results in [12] provide further evidence for the existence of driving behavior fluctuations. In contrast, fluctuations in longitudinal acceleration are manageable and enable more flexible speed trajectories [25].
For compatibility of flexibility and small behavior fluctuations, several studies design hybrid actions by discretizing parts of a continuous action space [26, 27], or using a parameterized action space [12, 25, 28], which generates both lateral discrete abstract targets and longitudinal continuous concrete acceleration commands. Additionally, [29] designs a dual-layer decision-making control model that combines parallel DQN and DDPG for hybrid output. [30] trains skill-agents for various driving objectives to output acceleration, from which DQN can flexibly select. However, the hybrid action methods mentioned above only generate discrete abstract targets, while our work further focuses on providing finer-grained abstract guidance for steering angle.
II-B Multi-Objective Policy Evaluation
@@
For AD problems, multiple attitudes should be considered, requiring the policy to be multi-objective compatible. The objectives are sometimes conflicting, like safety and driving efficiency [6]. The most common design is to linearly combine all attributes into a single, additive reward function for policy evaluation [31]. Specifically, the weights of this linearly expressed reward function are typically determined through manual design after multiple trial-and-error iterations [32], or by applying Inverse-RL to human demonstrations [33]. However, policy evaluation under this linear assumption may be inaccurate because the highest rewarding action may not be the one that enables multi-objective compatible driving, leading to reduced policy performance [34]. Additionally, a single critic representing multiple attitude rewards forces the learning of value coherence, which may not accurately reflect the true critic and degrade policy quality [35].
Some recent studies aim to develop architectures with multiple critics for multi-objective policy evaluation. In particular, they use several reward functions to separate key attributes from a single reward function, treating each attribute as an independent evaluation objective, such as safety[36], driving efficiency[37], and comfort[38]. Additionally,[39] introduces a pre-trained safe-critic to guide the policy towards safer actions. However, further investigation is required within a hybrid action-based RL framework.
II-C Policy Exploration Strategy
@@
Policy exploration helps agents discover potentially multi-objective compatible policies. A proper policy exploration strategy can significantly accelerate the learning process to converge to a viable policy faster [40]. However, the most common exploration strategy in RL is random exploration, such as $\epsilon$-greedy [41], which decreases exploration as training progresses. This randomized mechanism makes policy exploration lack of orientation and leads to repeated collection of experience samples, which reduces the training efficiency. Although some studies attempt to introduce noisy [42], reward novel state [43] or error of reward [44] to modify the exploration level, this does not change the nature of random exploration. Some other studies [45] attempt to use reward shaping to encourage exploration, but the manually imposed rewards heavily depend on the designer’s experience.
@@
Some studies [46, 47] use model ensemble technique to capture epistemic uncertainty and select actions that encourage the agent to explore high-uncertainty areas, thus accelerating policy training. Few studies leverage epistemic uncertainty in AD tasks to improve driving policy training efficiency [48]. In addition, most current methods only explore policies for a single action type.
@@ III-A Overall Framework
The method proposed in this paper is based on a hybrid parameterized action space for policy evaluation and improvement, considering multiple objectives to achieve multi-objective compatibility. Thus, the MDP can be rewritten as a new tuple $<{\cal S},{\cal H},\left[{{{\cal R}_{1}},\cdots,{{\cal R}_{N}}}\right],{\cal T},\gamma>$, where:
  • ${\cal H}$ represents the hybrid parameterized action space, where =${\cal H}=\left\{{\left({o,{a_{o}}}\right)\left|{{a_{o}}\in{{\cal A}_{\cal O}},{\rm{for\text{ }}}\forall o\in{\cal O}}\right.}\right\}$. The $o$ is the discrete action option selected from the discrete action option set ${\cal O}$. The $a_{o}$ can be seen as the continuous action parameter corresponding to $o$, drawn from the continuous interval ${\cal A}_{\cal O}$ corresponding to ${\cal O}$.
@@
As mentioned previously, AD tasks on structured roads require finer-grained abstract guidance to generate concrete control commands that reduce fluctuations while keeping flexibility in driving behavior. We design a hybrid action space that enables the agent to simultaneously output discrete actions $o$ and continuous action parameters $a_{o}$, ensuring optimality in both. These outputs are then used to generate both abstract guidance and concrete control commands. Specifically, lateral concrete control commands are generated by combining abstract guidance with prior knowledge, while longitudinal commands are directly derived from $a_{o}$.
@@
Figure 1: The overall framework of proposed HPA-MoEC. The actor of RL Method firstly generates the continuous action parameters $a_{o}$ based on states $s$, which are then input into the Multi-Objective Critics module along with $s$ for evaluating the value function. This module consists of $N$ ensemble-critics corresponding to the different attributes, and each of them is an ensemble of $M$ critics. The Exploration Strategy module then captures epistemic uncertainty from the ensemble-critics and selects the final hybrid action $(o,a_{o})$ that enhances training efficiency.
@@ III-B Policy and Value Function Representation
$Q\left({{s_{t}},{o_{t}},{a_{o,t}}}\right)=\mathbb{E}\left[{{r_{t}}+\gamma\mathop{\max}\limits_{o\in{\cal O}}\left\{{\mathop{\sup}\limits_{{a_{o}}\in{{\cal A}_{\cal O}}}{Q^{i}}\left({{s_{t+1}},o,{a_{o}}}\right)}\right\}}\right].$ (1)
HPA-MoEC consists of $N$ ensemble-critics, each composed of $M$ critics, resulting in a total of $N*M$ critics for value function evaluation. Specifically, each critic can estimate the value of the action $(o,a_{o})$ in state $s$ based on its focused attributes. Let ${Q_{ij}}$ represents the optimal value function evaluated by the $i$-th critic corresponding to ${\cal R}_{i}$:
$\begin{split}{Q_{ij}}({{s_{t}},{o_{t}},{a_{o,t}}})=\mathbb{E}\left[{r_{t}^{i}+\gamma\mathop{\max}\limits_{o\in{\cal O}}\left\{{\mathop{\sup}\limits_{{a_{o}}\in{{\cal A}_{\cal O}}}{Q_{ij}}\left({{s_{t+1}},o,{a_{o}}}\right)}\right\}}\right]\end{split}$ (2)
where $j\in[1,\cdots,M]$, $r_{t}^{i}={{\cal R}_{i}}(s,o,a_{o})$. However, finding the optimal continuous action $a_{o}$ is challenging in a hybrid parameterized action space. To overcome this, we assume that the value function is fixed, meaning that for any state $s$ and discrete action $o$, the $a_{o}$ depends on state $s$. At this stage, the problem of optimizing in the continuous space becomes determining the mapping from state $s$ to action $a_{o}$: ${\cal S}\to{{\cal A}_{\cal O}}$. By using a deterministic policy network $\mu(s;{\theta^{\mu}})$ to approximate this mapping, the continuous action $a_{o}$ can be obtained, with network parameters $\theta_{\mu}$. This policy network is known as the actor. Meanwhile, a value network is employed to approximate the value function $Q_{ij}$, with parameters $\theta_{ij}^{Q}$. Given the assumption that the value function is fixed, the MDP in the parameterized action space can be viewed as the process of exploring $\theta^{\mu}$ for a given $\theta_{ij}^{Q}$ :
@@
$\begin{split}{Q_{ij}}&\left({{s_{t}},{o_{t}},{a_{o,t}};{\theta_{ij}^{Q}}}\right)=\\ &\mathbb{E}\left[{r_{t}^{i}+\gamma\mathop{\max}\limits_{o\in{\cal O}}{Q_{ij}}\left({{s_{t+1}},o,\mu\left({{s_{t+1}};{\theta^{\mu}}}\right);\theta_{ij}^{Q}}\right)}\right].\end{split}$ (4)
The fundamental target for updating a single critic is:
$\begin{split}{y_{ij,t}}=r_{t}^{i}+\gamma\mathop{\max}\limits_{o\in{\cal O}}Q_{ij}^{\prime}\left({{s_{t+1}},o,\mu^{\prime}\left({{s_{t+1}};{\theta^{\mu^{\prime}}}}\right);\theta_{ij}^{Q^{\prime}}}\right),\end{split}$ (5)
@@
In our multi-objective policy evaluation architecture, each critic is not updated independently. For each ensemble-critic, every critic within shares the same driving attribute. Then, all ensemble-critics collaborate to guide the actor in learning a multi-objective compatible driving policy. To ensure consistency among all critics in evaluating driving behavior, the evaluation results —both for specific attributes and overall performance— should be fed back to each critic, for updating networks. Therefore, it is necessary to construct the critic’s update target from both the ensemble-critic perspective and the multi-objective compatible overall perspective. For the $i$-th ensemble-critic, its overall evaluation of the policy’s performance under a given attribute is the expectation of the value provided by the $M$ critics:
$\begin{split}{{\bar{Q}}_{i}}\left({{s_{t}},o,{a_{o,t}}}\right)&=\mathbb{E}{{}_{j\in\left[{1,\cdots,M}\right]}}\left[{{Q_{ij}}\left({{s_{t}},o,\mu\left({{s_{t+1}};{\theta^{\mu}}}\right);\theta_{ij}^{Q}}\right)}\right]\\ &=\frac{1}{M}\sum\limits_{j=1}^{M}{{Q_{ij}}\left({{s_{t}},o,\mu\left({{s_{t+1}};{\theta^{\mu}}}\right);\theta_{ij}^{Q}}\right)\left|{{}_{\forall o\in{\cal O}}}\right.}.\end{split}$ (6)
@@
$\begin{split}{{\bar{y}}_{i,t}}=r_{t}^{i}+\gamma\mathop{\max}\limits_{o\in{\cal O}}{{\bar{Q}}_{i}}^{\prime}\left({{s_{t+1}},o,{a_{o,t+1}}}\right),\end{split}$ (7)
@@
$\begin{split}{Q_{all}}\left({{s_{t}},o,{a_{o,t}}}\right)=\sum\limits_{i=1}^{N}{{\omega_{i}}{{\bar{Q}}_{i}}\left({{s_{t}},o,{a_{o,t}}}\right)}\left|{{}_{\forall o\in{\cal O}}}\right.,\end{split}$ (8)
@@
where $r^{all}$ combines the attribute rewards based on the attention level of each ensemble-critic, i.e., ${r^{all}}=\sum\nolimits_{i}^{N}{{\omega_{i}}{r^{i}}}$. The $Q_{all}^{\prime}$ is represented by the weighted sum of ${\bar{Q}_{i}}^{\prime}$, similar to Eq.(8).
@@
where ${{\bf{L}}_{t}}=\left[{{{\cal L}_{ij,t}},{{\cal L}_{i,t}},{{\cal L}_{all,t}},{{\cal L}_{conv,t}}}\right]$ represents the vector of loss function and ${{\bf{\lambda}}_{t}}=\left[{{\lambda_{1}},{\lambda_{2}},{\lambda_{3}},{\lambda_{4}}}\right]$ is the corresponding weight vector.

### v3

@@ Abstract
Reinforcement Learning (RL) has shown excellent performance in solving decision-making and control problems of autonomous driving, which is increasingly applied in diverse driving scenarios. However, driving is a multi-attribute problem, leading to challenges in achieving multi-objective compatibility for current RL methods, especially in both policy updating and policy execution. On the one hand, a single value evaluation network limits the policy updating in complex scenarios with coupled driving objectives. On the other hand, the common single-type action space structure limits driving flexibility or results in large behavior fluctuations during policy execution. To this end, we propose a Multi-objective Ensemble-Critic reinforcement learning method with Hybrid Parametrized Action for multi-objective compatible autonomous driving. Specifically, an advanced MORL architecture is constructed, in which the ensemble-critic focuses on different objectives through independent reward functions. The architecture integrates a hybrid parameterized action space structure, and the generated driving actions contain both abstract guidance that matches the hybrid road modality and concrete control commands. Additionally, an uncertainty-based exploration mechanism that supports hybrid actions is developed to learn multi-objective compatible policies more quickly. Experimental results demonstrate that, in both simulator-based and HighD dataset-based multi-lane highway scenarios, our method efficiently learns multi-objective compatible autonomous driving with respect to efficiency, action consistency, and safety.
@@ I Introduction
However, current RL methods still face several limitations in achieving compatibility with key driving objectives such as safety, efficiency, and action consistency [8, 9]. In particular, when addressing multi-attribute driving tasks, mainstream RL-based AD approaches exhibit shortcomings in both policy update and execution: (i) For policy updates, most rely on a single critic (value network) to evaluate and guide learning, making it difficult to efficiently explore multi-objective-compatible policies within a large and complex traffic state space; (ii) For policy execution, most approaches employ a single-type action space structure to handle hybrid road modality, which limits the policy’s ability to fully represent real driving behaviors and forces a trade-off among certain objectives.
In terms of policy update, employing a single critic (i.e., a single reward function) to evaluate policy performance fails to capture the strong coupling and potential conflicts among driving objectives. When multiple attributes of an AD task are combined into a single reward function, the agent may disproportionately focus on certain attributes during training [8]. As a result, some objectives may be neglected in specific states, leading to inaccurate value estimation and suboptimal policy performance. This can cause the agent to behave in ways that are misaligned with multi-objective expectations, such as becoming overly aggressive to maximize speed or excessively conservative to ensure safety. In contrast, multi-objective reinforcement learning (MORL) addresses this issue more effectively by constructing reward function vectors [10], enabling better compatibility among multiple objectives. Furthermore, complex traffic state spaces demand efficient exploration during policy updates. Most existing RL-based AD methods rely on random exploration, which prevents the agent from actively seeking out unknown regions and discovering potentially effective policies [11]. This random exploration often results in the collection of redundant experiences that contribute little to policy improvement, leading to inefficient convergence or even entrapment in local optima.
@@
For policy execution, using a single-type action space to generate either abstract or concrete driving behaviors constrains RL agents to discrete actions that lack flexibility or continuous actions that lack consistency. A common approach is to have the agent produce discrete, long-term driving goals, such as semantic decisions [12] or target points for path planning [13]. However, because the agent does not directly control the vehicle’s motion, its ability to adapt driving behavior flexibly is limited. While this long-term planning enhances action consistency, it reduces responsiveness to dynamic changes. Conversely, directly outputting short-term control commands [14] allows for greater flexibility, but often results in less consistent behavior, with frequent fluctuations and abrupt reactions to environmental changes.
@@
To alleviate the limitations of policy updating and execution in typical multi-objective AD tasks, this paper proposes a Multi-objective Ensemble-Critic reinforcement learning method with Hybrid Parametrized Action space (HPA-MoEC) for multi-objective compatibility. The HPA-MoEC adopts a MORL architecture focused on AD tasks. By defining multiple reward functions to decouple different driving attributes, each reward function guides an ensemble-critic to focus on specific driving objectives, thereby assisting the actor (policy network) in learning multi-objective compatible driving behaviors. The architecture further integrates a hybrid parameterized action space structure containing a discrete action set and its corresponding continuous parameters, which together generate driving actions that combine abstract guidance and concrete control commands. Additionally, uncertainty estimates from the ensemble-critic guide the agent to explore promising driving policies, facilitating more efficient exploration in unknown environments. Evaluation results on both simulation and HighD dataset-based multi-lane highway scenarios demonstrate that HPA-MoEC efficiently learns multi-objective compatible driving behaviors, significantly improving driving efficiency, action consistency, and safety. The main contributions are summarized as follows:
@@
    A MORL architecture compatible with multiple AD objectives is proposed, in which ensemble-critic focuses on a distinct objective using separate reward functions. Considering the safety-critical nature of AD, two driving objectives are defined and evaluated with two ensemble critics: one targeting overall performance, including interactivity, and the other dedicated to safety. By isolating the safety objective, the effectiveness of our architecture is demonstrated through improved safety performance in experimental results.
@@
    A hybrid parameterized action space structure is designed to combine finer-grained guidance and control commands to adapt to hybrid road modality. This hybrid action space consists of discrete actions and their corresponding continuous parameters, which together generate both abstract guidance and concrete control outputs. Our design achieves greater driving flexibility and reduced behavioral fluctuations, ensuring compatibility between driving efficiency and action consistency.
@@
    An epistemic uncertainty-based exploration mechanism is developed to enhance learning efficiency and complement the hybrid action space structure. By dynamically adjusting the direction and magnitude of exploration according to uncertainty and its trends, the agent is encouraged to more rapidly explore high-uncertainty regions for potentially effective policies. This exploration mechanism significantly improves the learning efficiency of multi-objective compatible policies.
@@ II Related Works
[…]
II-A Multi-Objective Policy Evaluation
@@
For AD problems, multiple attitudes should be considered, requiring the policy to be multi-objective compatible. The objectives are sometimes conflicting, like safety and driving efficiency [6]. The most common design is to linearly combine all attributes into a single, additive reward function for policy evaluation [16], typically based on mainstream RL algorithms such as Deep Q-Network (DQN) [17]and Soft Actor-Critic (SAC) [18]. Specifically, the weights of this linearly expressed reward function are typically determined through manual design after multiple trial-and-error iterations [19], or by applying Inverse-RL to human demonstrations [20]. However, policy evaluation under this linear assumption may be inaccurate because the highest rewarding action may not be the one that enables multi-objective compatible driving [10, 21], leading to reduced policy performance [22]. Additionally, a single critic representing multiple attitude rewards forces the learning of value coherence, which may not accurately reflect the true critic and degrade policy quality [23].
The MORL has recently attracted significant attention for its ability to address complex decision-making problems involving multiple, often conflicting, objectives [21]. A typical MORL approach employs an architecture with multiple critics to enable multi-objective compatible policy updates [10, 24, 25, 26]. Specifically, key attributes are separated from a single reward function by defining multiple reward functions, with each attribute treated as an independent evaluation objective [10]. Several AD-related studies have demonstrated the advantages of MORL in driving tasks by incorporating objectives such as safety [24], efficiency [27], and comfort [25]. Additionally, [26] introduces a pre-trained safe-critic to guide the policy towards safer actions. Building on these ideas, we propose HPA-MoEC, an advanced MORL architecture that integrates a novel hybrid action space and introduces epistemic uncertainty via ensemble-critic to enhance policy exploration and learning capability. Compared to previous MORL methods, HPA-MoEC achieves better and faster multi-objective learning in hybrid road modality.
II-B Action Space Structure
Many current RL-based AD methods use a single action type to control vehicle, which fail to be compatible with high driving flexibility and small behavior fluctuations. On one hand, some studies use a discrete action space to generate abstract behavior decisions, offering long-term targets that indirectly guide vehicle control. Specifically, [28, 29] use DQN and its improved versions to generate semantic lateral actions, such as left or right lane changes. Additionally, [30, 31] introduces longitudinal discrete acceleration and deceleration actions. To provide clearer guidance, some studies select from a discrete set of trajectories [32, 33] or directly generate the positions and desired speeds of target points [34]. However, these methods often reduce the alignment between agent outputs and driving behavior, as they rely on integration with a basic controller for vehicle control, which limits flexibility. On the other hand, some studies [35, 36] directly generate steering angles laterally and accelerations longitudinally from a continuous action space, aiming to enhance flexibility. However, fluctuations in the network’s output can cause frequent changes in steering angle and acceleration commands [37]. In scenarios with dedicated lanes, lateral fluctuations caused by steering angle variations will result in unpredictable paths. The experimental results in [14] provide further evidence for the existence of driving behavior fluctuations. In contrast, fluctuations in longitudinal acceleration are manageable and enable more flexible speed trajectories [38].
For compatibility of flexibility and small behavior fluctuations, several studies design hybrid actions by discretizing parts of a continuous action space [39, 40], or using a parameterized action space [14, 38, 41], which generates both lateral discrete abstract targets and longitudinal continuous concrete acceleration commands. Additionally, [42] designs a dual-layer decision-making control model that combines parallel DQN and Deep Deterministic Policy Gradient (DDPG) for hybrid output. [43] trains skill-agents for various driving objectives to output acceleration, from which DQN can flexibly select. However, the aforementioned studies fail to sufficiently integrate discrete and continuous actions, nor do they adequately account for the hybrid nature of road structures in complex driving environments. In contrast, our novel hybrid action space is specifically tailored to driving, providing abstract guidance compatible with hybrid road modality alongside continuous concrete control commands.
II-C Policy Exploration Mechanism
@@
Policy exploration helps agents discover potentially multi-objective compatible policies. A proper exploration mechanism can accelerate the learning process to converge to a viable policy faster [44]. However, the most common exploration strategy in RL is random exploration, such as $\epsilon$-greedy in DQN [45], random noise in TD3 [46], and the maximum entropy mechanism in SAC [18]. This randomized mechanism makes policy exploration lack of orientation and leads to repeated collection of experience samples, which reduces the training efficiency [47]. Although some studies attempt to introduce reward novel state [48] or error of reward [49] to modify the exploration level, this does not change the nature of random exploration. This inefficient exploration mechanism limits the potential performance of RL policies, particularly when pursuing multi-objective compatibility in complex traffic scenarios. Some other studies [50] attempt to use reward shaping to encourage exploration, but the manually imposed rewards heavily depend on the designer’s experience.
@@
Some studies [51, 52] use model ensemble technique to capture epistemic uncertainty and select actions that encourage the agent to explore high-uncertainty areas, thus accelerating policy training. Few studies leverage epistemic uncertainty in AD tasks to improve driving policy training efficiency [53]. Therefore, this paper develops an epistemic uncertainty-based exploration mechanism with multiple ensemble-critics for hybrid actions, enabling faster learning of multi-objective policy.
Figure 1: The overall framework of proposed HPA-MoEC. The actor of RL Method firstly generates the continuous action parameters $a_{o}$ based on states $s$, which are then input into the Multi-Objective Critics module along with $s$ for evaluating the value function. This module consists of NN ensemble-critics corresponding to the different attributes, and each of them is an ensemble of $M$ critics. The Exploration Strategy module then captures epistemic uncertainty from the ensemble-critics and selects the final hybrid action $(o,a_{o})$ that enhances training efficiency.
@@ III-A Overall Framework
The method proposed in this paper is based on a hybrid parameterized action space for policy evaluation and improvement, considering multiple objectives to achieve multi-objective compatibility. Thus, the MDP can be rewritten as a new tuple ${\cal S},{\cal H},\left[{{{\cal R}_{1}},\cdots,{{\cal R}_{N}}}\right],{\cal T},\gamma>$, where:
[…]
@@
To construct a fine-grained abstract guidance suitable for hybrid road modality, the designed hybrid action space enables the agent to simultaneously output discrete actions $o$ and continuous action parameters $a_{o}$, ensuring optimality in both. These outputs are then used to generate both abstract guidance and concrete control commands. Specifically, lateral concrete control commands are generated by combining abstract guidance with prior knowledge, while longitudinal commands are directly derived from $a_{o}$.
@@ III-B Policy and Value Function Representation
$Q\!\left({{s_{t}},{o_{t}},{a_{o,t}}}\right)\!=\!\mathbb{E}\!\left[{{r_{t}}\!+\!\gamma\mathop{\max\!}\limits_{\!o\in{\cal O}}\left\{\!{\mathop{\sup\!}\limits_{{a_{o}}\in{{\cal A}_{\cal O}}}\!{Q^{{}^{\prime}}}\!\left(\!{{s_{t+1}},o,{a_{o}}}\!\right)\!}\!\right\}}\!\right]\!.$ (1)
HPA-MoEC consists of $N$ ensemble-critics, each composed of $M$ critics, resulting in a total of NMN*M critics for value function evaluation. Specifically, each critic can estimate the value of the action $(o,a_{o})$ in state $s$ based on its focused attributes. Let $Q_{ij}$} represents the optimal value function evaluated by the $i$-th critic corresponding to ${\cal R}_{i}$:
$\begin{split}{Q_{ij}}(\!{{s_{t}},{o_{t}},{a_{o,t}}}\!)\!=\!\mathbb{E}\!\left[\!{r_{i,t}\!+\!\gamma\mathop{\max}\limits_{o\in{\cal O}}\!\left\{\!{\mathop{\sup}\limits_{{a_{o}}\in{{\cal A}_{\cal O}}}\!{Q_{ij}}\!\left({{s_{t+1}},o,{a_{o}}}\right)\!}\!\right\}}\!\right]\end{split}$ (2)
where $j\in[1,\cdots,M]$, $r_{i,t}={{\cal R}_{i}}(s,o,a_{o})$. However, finding the optimal continuous action $a_{o}$ is challenging in a hybrid parameterized action space. To overcome this, we assume that the value function is fixed, meaning that for any state $s$ and discrete action $o$, the $a_{o}$ depends on state $s$. At this stage, the problem of optimizing in the continuous space becomes determining the mapping from state $s$ to action $a_{o}$: ${\cal S}\to{{\cal A}_{\cal O}}$. By using a deterministic policy network $\mu(s;{\theta^{\mu}})$ to approximate this mapping, the continuous action $a_{o}$ can be obtained, with network parameters $\theta_{\mu}$. This policy network is known as the actor. Meanwhile, a value network is employed to approximate the value function $Q_{ij}$, with parameters $\theta_{ij}^{Q}$. Given the assumption that the value function is fixed, the MDP in the parameterized action space can be viewed as the process of exploring $\theta^{\mu}$ for a given $\theta_{ij}^{Q}$ :
@@
$\begin{split}{Q_{ij}}&\left({{s_{t}},{o_{t}},{a_{o,t}};{\theta_{ij}^{Q}}}\right)=\\ &\mathbb{E}\left[{r_{i,t}+\gamma\mathop{\max}\limits_{o\in{\cal O}}{Q_{ij}}\left({{s_{t+1}},o,\mu\left({{s_{t+1}};{\theta^{\mu}}}\right);\theta_{ij}^{Q}}\right)}\right].\end{split}$ (4)
To pursue higher returns, referring to the value network update target in DQN [17], the update target for a single critic is:
$\begin{split}{y_{ij,t}}=r_{i,t}+\gamma\mathop{\max}\limits_{o\in{\cal O}}Q_{ij}^{\prime}\left({{s_{t+1}},o,\mu^{\prime}\left({{s_{t+1}};{\theta^{\mu^{\prime}}}}\right);\theta_{ij}^{Q^{\prime}}}\right),\end{split}$ (5)
@@
In our architecture, each critic is not updated independently. For each ensemble-critic, every critic within shares the same driving attribute. Then, all ensemble-critics collaborate to guide the actor in learning a multi-objective compatible driving policy. To ensure consistency among all critics in evaluating driving behavior, the evaluation results—both for specific attributes and overall performance—should be fed back to each critic, for updating networks. Therefore, it is necessary to construct the critic’s update target from both the ensemble-critic perspective and the multi-objective compatible overall perspective. For the $i$-th ensemble-critic, its overall evaluation of the policy’s performance under a given attribute is the expectation of the value provided by the $M$ critics:
[…]
@@
$\begin{split}{{\bar{y}}_{i,t}}=r_{i,t}+\gamma\mathop{\max}\limits_{o\in{\cal O}}{{\bar{Q}}_{i}}^{\prime}\left({{s_{t+1}},o,{a_{o,t+1}}}\right),\end{split}$ (7)
@@
[…]
@@
where $r^{all}$ combines the attribute rewards based on the attention level of each ensemble-critic, i.e., ${r^{all}}=\sum\nolimits_{i}^{N}{{\omega_{i}}{r_{i}}}$. The $Q_{all}^{\prime}$ is denoted by the weighted sum of $\bar{Q}_{i}$$^{\prime}$, similar to Eq.(8).
@@
where ${{\bf{L}}_{t}}=\left[{{{\cal L}_{ij,t}},{{\cal L}_{i,t}},{{\cal L}_{all,t}},{{\cal L}_{conv,t}}}\right]$ represents the vector of loss function and ${{\bf{\lambda}}_{t}}=\left[{{\lambda_{1}},{\lambda_{2}},{\lambda_{3}},{\lambda_{4}}}\right]$ is the corresponding weight vector. By backpropagating the loss defined in Eq. 14, the value network $Q_{ij}$ can be updated iteratively.